OpenStudy (anonymous):

find dy/dx if y = (x^2 + 1)^3x

OpenStudy (anonymous):

write it as 3y=(x+(1/x)) now it looks easier i believe

OpenStudy (experimentx):

The other way!! \[ y = e^{\ln(x^2+1)*3x}\]

OpenStudy (anonymous):

what's that before the 3x?

OpenStudy (anonymous):

a plus?

OpenStudy (anonymous):

so tiny

OpenStudy (experimentx):

multiplication

OpenStudy (experimentx):

asterisk * <--- stands for multiplication

OpenStudy (anonymous):

im not sure what you did... you applied e to both sides?

OpenStudy (experimentx):

No ... just \( a = e^{\ln a}\) ... just simple rule of exponents and lograthims

OpenStudy (anonymous):

i just did chain rule so i guess that's not right

OpenStudy (experimentx):

???

OpenStudy (anonymous):

i got y' = 6x^2(x^2+1)^3x-1

OpenStudy (anonymous):

you should be getting y'=(x^2-1)/(3x^2)

OpenStudy (anonymous):

@experimentX I think your method is a lil bit complicated for a beginner in calculus

OpenStudy (experimentx):

I think either you or i misread question http://www.wolframalpha.com/input/?i=differentiate+%281%2Bx^2%29^%283x%29

OpenStudy (anonymous):

its power aa I misread the question my bad

OpenStudy (anonymous):

im lost.

OpenStudy (anonymous):

apply ln on both sides

OpenStudy (anonymous):

it becomes lny=(3x)(ln(x^2+1))

OpenStudy (anonymous):

now differentiate on both sides

OpenStudy (anonymous):

why do i have to apply ln

OpenStudy (anonymous):

it makes life easier

OpenStudy (anonymous):

because exponent is 3x?

OpenStudy (anonymous):

because there is an exponent

OpenStudy (anonymous):

if this was explicit differentiation then i could just use chain rule, correct?

OpenStudy (anonymous):

i got y' = 3xyln(x^2 + 1)

OpenStudy (experimentx):

put back the values of y..

OpenStudy (anonymous):

y' = 3x(x^2 + 1)^3x * ln(x^2 + 1)

OpenStudy (experimentx):

Take log on both sides \[ \ln y = 3x \ln(1+x^2)\] Differentiate both sides \[ \frac{1}{y} \frac{dy}{dx} = \frac{(6 x^2)}{(x^2+1)}+3 log(x^2+1)\] \[ y' = y \left ( \frac{(6 x^2)}{(x^2+1)}+3 log(x^2+1) \right )\] Put back the value of y, \[ y' = (1 + x^2)^{3x} \left ( \frac{(6 x^2)}{(x^2+1)}+3 log(x^2+1) \right )\] Answer!!!

OpenStudy (anonymous):

oh i forgot to differentiate right side

OpenStudy (anonymous):

log is the same as ln in this case?

OpenStudy (experimentx):

yeah it's ln

OpenStudy (anonymous):

ok! finally got it.